# Mosaic-IT: Free Compositional Data Augmentation Improves Instruction Tuning

### v2

@@ 4 Experimental Results
Table 2: The performance comparison on the MT-Bench, Alpaca Eval, and IF Eval Benchmarks. Rate(LC) in Alpaca Eval represents the length-controlled win rates. In IF Eval, Prompt, and Inst represent Prompt-level and Instruction-level accuracy; S and L represent Strict and Loose versions.
Model Dataset Method MT-Bench $\uparrow$ Alpaca Eval 2 $\uparrow$ IF Eval $\uparrow$
1-round 2-round Rate (LC) Rate Prompt (S) Inst (S) Prompt (L) Inst (L)
Mistral 7B Alpaca-GPT4 Baseline 6.44 5.26 3.98 7.28 32.53 42.93 35.86 45.92
Mosaic-IT 7.11 4.69 5.00 7.81 37.15 48.56 38.08 50.23
Wizard-70k Baseline 6.21 4.70 4.13 6.46 39.56 49.88 41.96 53.00
Mosaic-IT 6.95 4.32 4.44 7.56 40.85 51.80 45.47 56.47
@@
Llama-3-8B Vicuna Baseline 1.000 52.51 2.15 1.36 6.70 5.06 21.26 33.45
Mosaic-IT 1.234 55.62 3.09 2.05 6.85 5.40 31.42 45.56
Magpie Baseline 1.000 56.15 9.22 13.74 8.10 7.08 35.67 47.72
@@
Phi-3 Vicuna Baseline 1.000 62.06 4.20 2.74 5.34 4.18 30.50 43.17
Mosaic-IT 1.083 62.30 5.95 3.83 5.89 4.53 32.35 41.85
Magpie Baseline 1.000 62.90 13.82 17.68 7.78 6.42 44.36 55.52
@@
Gemma2-2B Vicuna Baseline 1.000 48.90 1.72 1.31 6.69 5.25 23.66 35.61
Mosaic-IT 1.266 51.31 1.90 1.38 6.93 5.26 24.03 36.93
Magpie Baseline 1.000 46.37 5.35 7.77 4.57 3.23 21.81 32.49
@@ 4.1 Main Results
In this section, we present the evaluation results comparing our methods with the baseline methods on several baseline models (Mistral-7B Jiang et al. (2023), Llama2-7B Touvron et al. (2023b), Llama2-13B) and instruction tuning datasets (Alpaca-GPT4 Peng et al. (2023), Alpaca Taori et al. (2023), WizardLM-70k Xu et al. (2023)), on Two general evaluation settings (Pair-Wise Comparison and Open LLM leaderboard) described above, as shown in the Table 1. Pair-wise Winning Score indicates the result directly comparing our models with the corresponding baseline models, which is calculated as (Num(Win)--Num(Lose))///Num(All) $+1$. These values that are greater than $1.0$ represent better responses generated by our models. The performances on the Huggingface Open LLM Leaderboard are also presented, and we bold the greater average values for each comparison. The consistent outperforming results on different base models and datasets represent the effectiveness and robustness of our methods.
@@
To better understand how our method improves the instruction-following abilities of LLMs, we further compare the performance on other Three benchmarks for fine-grained analysis based on the Mistral-7B base model with two datasets as shown in Table 2. On the MT-Bench, the 1-round scores of our method are higher, indicating that our method mainly improves the response quality for single-round conversations, which is reasonable as the meta instructions only focus on single-round formats. On the Alpaca Eval benchmark, our method has a consistent improvement with or without the Length Control (LC), indicating that the improvement of response qualities does not directly originate from the length of responses. On the IF Eval benchmark, our method consistently improves the performances on all 4 different settings, both Prompt-level and Instruction-level, both Strict version and Loose version. Compared with the previous benchmarks, IF Eval mainly focuses on the constraint-following ability of LLMs. The consistent improvement in this benchmark represents that our method not only improves the response qualities of the LLMs but also improves their controllability regarding formats. Given that our method is a cost-free augmentation technique that does not rely on any additional models, the observed improvements are remarkable.
Moreover, to further verify the effectiveness of our method, more experiments on different model families and data families are conducted, as shown in Table 3, including Llama-3-8B Dubey et al. (2024), Phi-3 Abdin et al. (2024), and Gemma2-2B Team et al. (2024) models on Vicuna 1M Zheng et al. (2024a), and Magpie Xu et al. (2024b) datasets. For these two datasets, $300$k data are randomly sampled to verify the scalability of our method when dealing with large amounts of instruction-tuning data. The performances of our models consistently outperform the baseline models across different model families and data sources, ranging from diverse data qualities.
@@
Further Human Evaluations are conducted on Mistral-7B with Alpaca-GPT4 and WizardLM dataset. For the comparison on (1) Alpaca-GPT4: the model using Mosaic-IT wins on $68$ out of $100$ instruction, ties on $3$, and losses on $29$ instructions; on (2) WizardLM: the model using Mosaic-IT wins on $63$ out of $100$ instruction, ties on $6$, and losses on $31$ instructions. This human evaluation also further verifies the effectiveness of our Mosaic-IT.

### v3

@@ 3 Experimental Results
Llama-3-8B Magpie Baseline 1.000 56.15 9.22 13.74 8.10 7.08 35.67 47.72
@@
Phi-3 Magpie Baseline 1.000 62.90 13.82 17.68 7.78 6.42 44.36 55.52
@@
Gemma2-2B Magpie Baseline 1.000 46.37 5.35 7.77 4.57 3.23 21.81 32.49
@@
Mistral 7B Alpaca-GPT4 Baseline 1.000 59.70 3.98 7.28 6.44 5.26 35.86 45.92
Mosaic-IT 1.349 63.65 5.00 7.81 7.11 4.69 38.08 50.23
Wizard-70k Baseline 1.000 57.86 4.13 6.46 6.21 4.70 41.96 53.00
Mosaic-IT 1.161 61.11 4.44 7.56 6.95 4.32 45.47 56.47
Table 2: The performance comparison across multiple model families and datasets on five evaluation metrics. Rate(LC) in Alpaca Eval represents length-controlled win rates. In IF Eval, P(L) and I(L) represent Prompt-level and Instruction-level accuracy in the Loose setting.
@@ 3.1 Main Results
In this section, we present the evaluation results comparing our methods with the baseline methods on 6 baseline models (Mistral-7B Jiang et al. (2023), Llama2-7B Touvron et al. (2023b), Llama2-13B, Llama-3-8B Dubey et al. (2024), Phi-3 Abdin et al. (2024), Gemma2-2B Team et al. (2024)) and 4 instruction tuning datasets (Alpaca-GPT4 Peng et al. (2023), Alpaca Taori et al. (2023), WizardLM-70k Xu et al. (2023)), Magpie Xu et al. (2024b), on 5 commonly used evaluation metrics and additional Human Evaluation. Detailed experimental setup and descriptions of evaluation metrics can be found in Appendix B.
@@
Table 1 shows the results on 2 general evaluation settings (Pair-Wise Comparison and Open LLM leaderboard). Pair-wise Winning Score indicates the result directly comparing our models with the corresponding baseline models, which is calculated as (Num(Win)--Num(Lose))///Num(All) $+1$. These values that are greater than $1.0$ represent better responses generated by our models. The performances on the Huggingface Open LLM Leaderboard are also presented, and we bold the greater average values for each comparison. The consistently outperforming results on different base models and datasets represent the effectiveness and robustness of our methods. Results on more advanced baseline models and datasets can be found in Table 2. Our method shows consistent improvements compared with baseline models. The performance gains in pair-wise comparison indicate our method helps LLMs generate more detailed and accurate responses, and the performance gains in the open leaderboard indicate our method helps alleviate harm to the intrinsic capabilities of base LLMs.
@@
To better understand how our method improves the instruction-following abilities of LLMs, we further compare the performance on the other 3 benchmarks for fine-grained analysis as shown in Table 2. On the Alpaca Eval 2 benchmark, our method has a consistent improvement with or without the Length Control (LC), indicating that the improvement of response qualities does not directly originate from the length of responses. On the MT-Bench, the 1-round scores of our method are consistently higher, while the 2-round scores slightly fluctuate, indicating that our method mainly improves the response quality for single-round conversations since our meta instructions only focus on single-round scenarios in this version. On the IF Eval benchmark, our method consistently improves the performance on different settings, both Prompt-level and Instruction-level. Compared with the previous benchmarks, IF Eval mainly focuses on the constraint-following ability of LLMs. The consistent improvement in this benchmark represents that our method not only improves the response qualities of the LLMs but also improves their controllability regarding formats.
@@
Further Human Evaluations are conducted on Mistral-7B with the Alpaca-GPT4 and WizardLM datasets. For the comparison on (1) Alpaca-GPT4: the model using Mosaic-IT wins on $68$ out of $100$ instruction, ties on $3$, and losses on $29$ instructions; on (2) WizardLM: the model using Mosaic-IT wins on $63$ out of $100$ instruction, ties on $6$, and losses on $31$ instructions. This human evaluation also further verifies the effectiveness of our Mosaic-IT.
To conclude, our Mosaic-IT shows consistent improvement in instruction-following and constraint-following ability and response quality with a reduction of approximately $80\%$ of the training time cost. Given that our method is a cost-free data synthesis technique that does not rely on any additional human/LLM generation, the observed improvements are remarkable.
Time (min) 827 121 129 133 133 143 716 426 305 245 202 173
@@ 3.2 Ablation Studies
Detailed ablations are in Appendix C.
Ablation on Mosaic Strategies investigates how different mosaic strategies affect LLM performances. We find out that further implementing Advanced Strategies (Format, Permute, Maskout) improves LLM performance as they largely diversify and complicate the instructions, only implementing the Primary Strategy. Ablation on the Max Number of Instructions investigates the trend between the number of instructions that are concatenated together and the LLMs’ performances. We find out that in the uniform distribution setting, more of the instructions are concatenated together, as this process makes the concatenated instruction complex. Ablation on the Distribution of Number of Instructions investigates how different distributions affect the LLMs’ performance. It is revealed that, except for the maximum number of instructions concatenated together, the distribution is also important. Ablation on Semantic Grouping investigates how semantic grouping, i.e., grouping instructions with similar semantic meaning rather than pure random grouping, affects performance. We show that semantic grouping has its own benefits, while random selection is much more convenient.
4 Further Discussion
4.1 Improving Efficiency
One of the benefits of our method is the efficiency of the training process. Given an existing dataset, our mosaic processes largely decrease the number of total overall instructions and the total number of gradient descents, leading to a reduction in the training process. The detailed comparison is shown in Table 3, which is based on the Mistral-7B model on the Alpaca-GPT4 dataset. The time is calculated based on four NVIDIA A100 Graphic Cards. As shown, our method greatly decreases the training time to approximately $16\%$ to $25\%$ while achieving better performances, especially when there are mosaic samples with larger permutation counts. Please note that the time reduction is ratio-based, when larger datasets or models are trained, the absolute time reduction gap between baseline methods and our methods will be much more obvious.
4.2 Alleviating Memorizing
In the original instruction tuning process, each data sample will be trained several times for LLMs without changes to the instructions and responses. This training process poses risks to the potential memorizing effects on training samples, which can be partially indicated by the “stair-like” training loss curves as shown in Figure 3. In the figure, all the training settings are kept the same between the baseline models and Mosaic-IT models, including the Learning Rate, Warm-up Ratio, Learning Rate Schedule (Cosine), Batch Size, etc. For the baseline methods, the training loss hardly decreases within each epoch of training but drops dramatically when the LLMs meet the same training samples again, which indicates a potential memorizing effect of training samples and potential overfitting. However, when utilizing our method, the random mosaics of original instructions with diverse and complex meta-instructions largely diversify the overall training instructions. Although each original data sample will still be seen by LLMs several times during training, the overall context varies dramatically as each original sample is only an atomic element of the overall mosaic sample, indicating that there will be no identical overall instructions during the whole training process. Thus, this augmentation largely alleviates the potential memorizing and overfitting problems as shown in the figure, where the training loss decreases smoothly, representing the gradual learning process.
GPT-3.5-turbo (Sequential) 0.357 0.014 0.336 0.055 0.303 0.064
GPT-3.5-turbo (Random) 0.315 0.124 0.330 0.156 0.198 0.312
GPT-4-turbo (Sequential) 0.176 0.000 0.137 0.000 0.140 0.000
GPT-4-turbo (Random) 0.139 0.000 0.153 0.014 0.101 0.005
5 Why It Works
5.1 Preliminaries: Performance Degradation
5.2 Further Analysis
Mosaic-IT trains LLMs to follow meta-instructions for compositional reasoning.
Previous methods train LLMs to produce a response for a single instruction or query. Instead, our method produces a compositional data synthesis method to train LLMs to generate multiple responses for multiple instructions in diverse forms (e.g., order, mask, format) specified by different meta-instructions. It also enforces LLMs to partition the input context correctly and manage the interference and dependencies among multiple instructions. These are critical to developing and improving the compositional reasoning capabilities of LLMs, which have not been covered by mainstream instruction-tuning frameworks.
Mosaic-IT creates more challenging and complex instructions to further improve LLMs’ instruction-following capabilities.
Mosaic-IT’s composition of multiple instructions and the diverse meta-instructions create more challenging and complex instruction-tuning data for LLMs. Moreover, since we do not rely on data synthesis using LLMs but solely apply some rules to existing data, the correctness and quality of the augmented data are guaranteed. As shown in Table 4, even powerful LLMs like GPT4 can not follow concatenated instructions. Besides, it has been widely accepted that such challenging and complex instructions improve LLMs’ instruction-following capability Zhao et al. (2024); Wu et al. (2024); Ding et al. (2023); Li et al. (2023a); Liu et al. (2023a); Li et al. (2024c, b); Guo et al. (2024). Mosaic-IT follows this intuition by making the instruction more challenging and complex in order to improve LLMs. Different from previous methods relying on humans or stronger teacher LLMs to create the challenging samples, Mosaic-IT does not require any humans/models to create the augmentations.
To quantitatively evaluate the difficulty and complexity of instruction-tuning data, InsTag Lu et al. (2023) proposes a ChatGPT-based method (Number of InsTag), while Cherry Li et al. (2024e) proposes a perplexity-based Instrutcion-Following Difficulty (IFD) score. We compute these two metrics on the Alpaca and WizardLM70k datasets to verify the effectiveness of our method in improving the difficulty/complexity:
Number of InsTag: The number of InsTag is used to measure the complexity of the instructions. A larger value of the Number of InsTag indicates the intentions of the instruction are complex and benefit the LLM instruction tuning process. For the experiments below, we prompt GPT4o with the exact prompt provided in the paper to generate the Instags. As shown in Table 5, Mosaic-IT largely increases the average number of InsTag, indicating a large increase in instruction intention complexity, further leading to better performance.
IFD score: IFD score is a perplexity-based metric used to evaluate the instruction-following difficulty of a given instruction-response pair. A higher IFD score indicates that it is hard for the current model to build a connection between the instruction and the corresponding response, so it can be used to select training data beneficial for LLM instruction tuning. For the experiments below, we utilized the IFD score computed on GPT2. As shown in Table 5, Mosaic-IT increases IFD scores, indicating an increase in the instruction-following difficulty, which leads to an improvement in performance.
Method Ins Tag IFD
Alpaca Wizard-70k Alpaca Wizard-70k
Original 2.62 4.20 0.60 0.67
Mosaic-IT 9.75 10.93 0.76 0.79
Table 5: The comparison between the original dataset and Mosaic-IT enhanced dataset with respect to the Number of InsTag and the IFD score.
5.3 Qualitative Example
To better understand the differences between the responses generated by the baseline model and the Mosaic-IT model, a pair of qualitative examples are presented in Figure 4 (baseline model) and Figure 5 (Mosaic-IT model). The example instruction is selected from the WizardLM test sets and the models are Mistral-7B trained on Alpaca-GPT4 with or without our method.
Some of the real-world instructions can be complex and hard to answer all at once but require LLMs to “decompose” the original overall instruction into pieces to respond. Due to the lack of this kind of difficult data in the training set, the capability of most LLMs is largely restricted as shown in Figure 4. The baseline LLM directly ignores the “sub-query” of explaining “protocols and standards such as TCP/IP, HTTP, FTP, DNS, DHCP, and ARP”. On the contrary, as a cost-free data synthesis method, Mosaic-IT simulates this kind of instruction by using original easy instructions, thus equipping LLMs with the capability to respond to these difficult instructions without ignoring some parts of it, as shown in Figure 5.
The better performances on Alpaca Eval 2 (LC) show that the performance gains of Mosaic-IT do not purely originate from longer responses. On the contrary, Mosaic-IT actually makes the lengths of training data more diverse compared with the original data, e.g., some of the samples are the concatenation of only 1 original sample, while some of them are the concatenation of 10 samples. This ensures diverse generation lengths according to the given instructions rather than being dominated by the original length distribution.
6 Conclusion
We introduce Mosaic Instruction Tuning (Mosaic-IT), a novel, human/model-free method to enhance instruction tuning for LLMs. By concatenating multiple instruction-response samples and using higher-level meta-instructions, Mosaic-IT improves multi-step and format-following capabilities. Our evaluations show superior performance and an $80\%$ reduction in training costs compared to the original methods. Mosaic-IT’s simplicity and efficiency make it a scalable solution for improving LLMs without extensive human intervention or resource-intensive teacher models. Our results highlight the potential of innovative data synthesis techniques in advancing LLM capabilities.
Limitations
Appendix A Qualitative Examples
To better understand the differences between the responses generated by the baseline model and the Mosaic-IT model, a pair of qualitative examples are presented in Figure 4 (baseline model) and Figure 5 (Mosaic-IT model).
  Identify some of the main components of a computer network such as hosts, routers, switches, hubs, links, and interfaces. Explain how they communicate and coordinate with each other using protocols and standards such as TCP/IP, HTTP, FTP, DNS, DHCP, and ARP. Describe how data is transmitted and received over a network using concepts such as packets, frames, headers, addresses, ports, sockets, and checksums.
A computer network is composed of several components that work together to facilitate communication and data transfer between devices. These components include:
1. Hosts: A host is a device that is connected to the network, such as a computer, smartphone, or printer. Each host has a unique IP address, which is used to identify it on the network.
2. Routers: A router is a device that connects multiple networks together and directs traffic between them. It uses protocols such as IP to route data packets to their destination.
3. Switches: A switch is a device that connects multiple devices together on a single network. It uses MAC addresses to direct data packets to their destination.
4. Hubs: A hub is a device that connects multiple devices together on a single network. It broadcasts data packets to all connected devices, which allows them to communicate with each other.
5. Links: A link is the physical connection between two devices on a network. It can be a cable, a wireless connection, or a combination of both.
6. Interfaces: An interface is the part of a device that connects it to the network. It can be a network card, a USB port, or a wireless card.
All of these components communicate and coordinate with each other using protocols and standards such as TCP/IP, HTTP, FTP, DNS, DHCP, and ARP. These protocols define how data is transmitted and received over a network, how devices are addressed, and how traffic is routed.
Data is transmitted over a network in packets or frames. Each packet or frame contains a header, which includes information such as the destination IP address, the source IP address, and the type of data being transmitted. The data is then transmitted over the network, and the receiving device uses the information in the header to identify the data and direct it to the appropriate application or program.
In summary, a computer network is composed of several components that work together to facilitate communication and data transfer. These components communicate and coordinate with each other using protocols and standards, and data is transmitted over the network in packets or frames, using headers to identify the destination, source, and type of data.
Figure 4: An example from the WizardLM test set with the response generated by the Baseline model.
A computer network is made up of several components that work together to facilitate communication and data transfer between devices. Some of the main components include:
1. Hosts: These are the devices that are connected to the network and use its resources, such as computers, printers, and servers.
2. Routers: Routers are devices that direct traffic between networks, and are used to connect multiple networks together.
3. Switches: Switches are used to direct traffic within a network, and are used to connect multiple devices together on a single network.
4. Hubs: Hubs are devices that connect multiple devices together on a single network, and are used to distribute data to multiple devices at once.
5. Links: Links are the physical connections between devices on a network, such as Ethernet cables or wireless connections.
6. Interfaces: Interfaces are the points of connection between the network and the devices on it, such as network cards or wireless adapters.
For these components to communicate and coordinate with each other, they use a variety of protocols and standards. Protocols are sets of rules that dictate how data is transmitted and received, while standards are agreed-upon methods or formats that allow for interoperability between different systems. Some common protocols and standards used in computer networking include:
- TCP/IP: This is a suite of protocols that is used to communicate between devices on a network. It includes protocols for addressing, routing, and transporting data.
- HTTP: This is a protocol used for accessing and transferring data over the web. It is used to communicate between web servers and clients.
- FTP: This is a protocol used for transferring files over a network. It is used to communicate between FTP servers and clients.
- DNS: This is a protocol used for translating domain names to IP addresses. It is used to coordinate communication between devices on a network.
- DHCP: This is a protocol used for assigning IP addresses to devices on a network. It is used to coordinate communication between devices on a network.
- ARP: This is a protocol used for finding the MAC address of a device on a network. It is used to coordinate communication between devices on a network.
When data is transmitted over a network, it is broken down into small packets, each of which contains information about where it needs to go and what it contains. These packets are sent over the network, hopping from one device to the next until they reach their destination. At each hop, the device examines the packet’s header to see where it needs to go next, and then forwards it along the appropriate link.
Once the packet reaches its destination, the receiving device examines the header to see if it is the intended recipient. If it is, the device will extract the data from the packet and use it, otherwise, the packet will be discarded.
To ensure that the data in the packets is not corrupted during transmission, checksums are used. A checksum is a mathematical calculation performed on the data in the packet, which produces a value that can be used to verify that the data has not been altered during transmission. If the checksum of the data in a packet is incorrect when it arrives at its destination, it is discarded, and the sender is notified of the error.
In summary, data is transmitted over a network by being broken down into small packets, each of which contains information about where it needs to go and what it contains. These packets are sent hopping from one device to the next, using a variety of protocols and standards to coordinate their transmission and reception. Checksums are used to ensure that the data in the packets has not been corrupted during transmission, and any errors are detected and corrected.
Figure 5: An example from the WizardLM test set with the response generated by the Mosaic-IT model.
Appendix B Experimental Setup
B.1 Implementation Details
For the three base pre-trained models, Llama2-7B, Llama2-13B Touvron et al. (2023b), and Mistral-7B Jiang et al. (2023), we utilize the prompt and code base from Vicuna Chiang et al. (2023) and flash attention Dao et al. (2022). The overall training arguments are aligned with the common training configuration. The Adam optimizer Kingma and Ba (2017) is utilized with the batch size of $128$ and with the max token length of $2048$. When training the baseline models Llama2-7B and Llama2-13B, the maximum learning rate is set to $2\times 10^{-5}$ with the warmup rate as $0.03$ for $3$ epochs. When training the baseline models on Mistral-7B, the maximum learning rate is set to $1\times 10^{-5}$ with the warmup rate as $0.1$ for $3$ epochs. For the three models, Llama-3-8B Dubey et al. (2024), Phi-3 Abdin et al. (2024), and Gemma2-2B Team et al. (2024), we utilize the code base from LLaMA-Factory Zheng et al. (2024b). The max token length is set with $4096$ following the modern settings and we train the model for $2$ epochs. Other parameters are kept the same as the above.
B.2 Training Dataset
The Alpaca dataset Taori et al. (2023) comprises $52,000$ instruction-following samples and is constructed utilizing the self-instruct paradigm Wang et al. (2023b). This dataset was produced by employing OpenAI’s text-davinci-003 model. Characterized as a classical dataset with moderate quality attributes, the Alpaca dataset serves as an initial platform to validate our methodology. To further substantiate our approach using a dataset of inherently high quality, we also applied our method to the Alpaca-GPT4 dataset Peng et al. (2023), which features responses generated by GPT4. The WizardLM dataset Xu et al. (2023) is also utilized in our method, which contains $70,000$ samples created by the evolution algorithm proposed by them. With ChatGPT-3.5 utilized, the data quality on WizardLM is largely guaranteed. The Vicuna 1M dataset Zheng et al. (2024a) is a large-scale dataset containing one million real-world conversations with 25 state-of-the-art LLMs. Due to the computation budget, $300$k instances are randomly sampled for our experiments. Magpie dataset Xu et al. (2024b) is the most recent SOTA synthetic dataset with $300$k samples.
B.3 Evaluation Metrics
Pair-wise Comparison by using powerful LLMs like GPT-4 is recently widely accepted and becoming a common practice Touvron et al. (2023b); Chiang et al. (2023); Dettmers et al. (2023); Liu et al. (2023b); Chiang and Lee (2023). The evaluation of responses from LLMs, especially in open-domain contexts where definitive ground truth is hard to establish, continues to be an intricate and evolving research domain. Recent studies, however, have indicated a notable alignment between GPT-4’s performance evaluations and human assessments Zheng et al. (2023); Li et al. (2023c); Sottana et al. (2023), thereby establishing a credible foundation for this evaluative methodology. We adopted test instruction sets from WizardLM Xu et al. (2023), comprising 218 diverse, human-curated instructions for pair-wise comparison. We directly follow the evaluation framework proposed by Chen et al. (2023); Li et al. (2024e), which evaluates responses on a scale spanning from 1 to 10 across multiple dimensions. To further address positional bias, as discussed by Ko et al. (2020); Wang et al. (2023a), the comparison is conducted in two distinct sequences, LLM1’s response first and then LLM2’s response first, ensuring a fair assessment of model performance. Evaluation outcomes are categorized into ’win-tie-loss’ for each instruction. The detailed evaluation prompt is provided in Appendix H.
Appendix C Detailed Ablation Studies